# Significant impact of Galactic dark matter particles on annihilation signals from Sagittarius analogues

### v1

@@ 5 Gamma-ray data and model fitting
We consider a $15^{\circ}\times 15^{\circ}$ ROI centered on M54/Sgr. For our likelihood maximization, we take a ${}^{\circ}$ angular pixelation and use the MINUIT optimizer method within gtlike. We use an input source model that includes all sources in the 4FGL-DR3 catalog [12] out to a square of $20^{\circ}\times 20^{\circ}$. Including sources beyond the ROI ensures that sources on the edge of the ROI are properly modeled. For the interstellar emission model we use the recommended gll_iem_v07.fits, and for the isotropic emission we use iso_P8R3_SOURCE_V3_V1.txt. To report the significance of our sources, we report the TS value obtained upon running gta.sed, which takes the source following the procedure outlined above and then performs a bin-by-bin likelihood fit assuming a power law spectralization per bin with an index set to $\Gamma=-2$. Finally, for all TS values reported we assume a DM annihilation spectrum, as defined by DMFitFunction provided in FermiPy assuming annihilation via the $b\bar{b}$ channel. For this spectrum, we fit the $\mathcal{J}$-factor to the total integrated $\mathcal{J}$-factor for a given source map, fix the mass to the best-fitting mass found in ref. [13] of 29.6 GeV, and allow $\langle\sigma v\rangle$ to float.
@@
The flux of annihilating DM from the MW halo dominates the Sgr source for p and d-waves ($\mathcal{J}_{\text{Sgr}}/\mathcal{J}_{\text{MW}}\sim 0.0001-0.1\%$, where $\mathcal{J}_{\text{MW}}$ denotes the MW $\mathcal{J}$-factor integrated over the area of the target), and is similar to the source flux in the central regions for the s-wave ($\mathcal{J}_{\text{Sgr}}/\mathcal{J}_{\text{MW}}\sim 0.1-10\%$). For the Sommerfeld case, the Sgr source flux is always larger than the MW halo emission ($\mathcal{J}_{\text{Sgr}}/\mathcal{J}_{\text{MW}}\sim 5-20\%$). When fitting both the Sgr source and the MW diffuse over the entire ROI, we find an overall poor fit quality, because we are modeling the diffuse emission over such a large region of sky. This likely means that an appropriate handling of the MW DM emission requires unique templates for the diffuse MW emission arising from cosmic ray interactions with gas and the interstellar radiation field, which is beyond the scope of this work. Therefore, for the analysis in this section we model only the emission associated with the Sgr source. In this way, we can compare qualities of the source independent of the MW diffuse emission, such as the effect of including the unbound DM particles in our templates.
@@
To model M54/Sgr in gamma-rays based on our simulations results, using the results from figure 4, we create a flux map from our pixel-by-pixel calculation of the $\mathcal{J}$-factor. Since the simulated Sgr source is not smooth, we smooth the counts maps obtained from the simulation with a Gaussian kernel of size ${}^{\circ}$, ensuring that the total $\mathcal{J}$-factor for each map is preserved. We find that this smoothing step is necessary to obtain a good model of the source, which is generally best fit by a point source model, or a smooth model with a small, $\lesssim$ 1 degree extension [13]. Since Sommerfeld sources are the most promising candidates for detection, we focus on these models for the remainder of this section. For this portion of the analysis, we focus on the results of the Sgr analogue with the smallest angular size, Sgr 25 (angular size $\sim 1.4^{\circ}$), and compute TS values for the Sommerfeld model. Choosing the Sgr analogue with the smallest angular size ensures the best fit to the source given its small extension, as attempting to fit to the other larger analogues produces poor overall fit qualities within the ROI. For the Sommerfeld model of Sgr 25, we find a TS $\sim 5$. This includes both bound and unbound DM particles, as described above. For a map which includes only the bound DM particles to the Sgr analogue, we find a TS of $\sim 14$.
@@ 6 Discussion and conclusions
Using analogues of the Sgr dSph identified in the Auriga simulations, we have computed $\mathcal{J}$-factors for general models in which the annihilation cross section is velocity-dependent. We provide predictions for both the DM component that is associated with the Sgr dSph, and with the MW foreground component along the line-of-sight to the dSph. For s-wave and Sommerfeld models, we find that emission from the central source is generally larger than that from the foreground MW component in the central region ($\lesssim 0.5^{\circ}$) and therefore would most likely be identifiable as a point source in the data. We further find that the $\mathcal{J}$-factors that we compute from the simulations are in agreement with those obtained from equilibrium models presented in ref. [2], implying that the latter are accurate for the central core region of Sgr.
@@
For velocity-dependent models, for the first time we establish the existence of a correction factor due to unbound DM particles from the MW that are spatially overlapping with Sgr. These unbound particles have a relative velocity distribution characteristic of the smooth DM component of the MW at the radius of the Sgr analogue. The correction factor results from annihilation of these unbound DM particles with themselves and with DM particles bound to the Sgr analogues. For p-wave models, we find that this correction varies from $\sim 2-20$, depending on the exact Sgr analogue used. For d-wave models, we find a larger correction varying between $\sim 2-200$. This correction is different than the more traditional “boost” factor that has been discussed often in the literature, which results from DM substructure [3, 46, 47]. On the other hand, for Sommerfeld models, the correction reduces the $\mathcal{J}$-factor by up to an order of magnitude. This implies that the sensitivity to the annihilation cross section for Sommerfeld models is likely weaker than obtained in previous analyses that used only models for the bound component of the system [19].
@@
Due to the angular extension of our models of $\sim 1-3$ degrees, we find that our Sgr templates generally do not provide as good a fit to the gamma-ray data as a simpler point source model. For example, examining in detail the Sommerfeld model for the case of our Sgr analogue with the smallest angular extent (Sgr 25, with an angular extent of ${}^{\circ}$), on Fermi-LAT data, we find a TS $\sim 5$ for the template including emission from bound and unbound particles. Restricting to only the particles identified as bound to the Sgr analogue, the TS increases to $\sim 14$, which is expected given the increase in the $\mathcal{J}$-factor for the latter case. We note that these TS values are less significant than the TS values for the point-like scenario for the M54/Sgr source as reported in refs. [12] and [13]. More detailed modeling of combined emission from the MW halo and Sgr analogue DM annihilation would require specialized detailed diffuse model templates for the region, which is beyond the scope of this work.
@@
We finally note that it is likely that the correction factor that we derive is not restricted to just Sgr-like systems. In fact, it should exist for all subhalos that are within close enough proximity to the Galactic center where there is a significant DM density that is spatially coincident with the subhalo or for subhalos that are in an overdense region in the Galaxy. Furthermore, a large difference between the average relative velocity of the DM particles bound to the subhalo and to the smooth MW halo in that region also contributes to this correction factor. Particularly interesting systems to investigate in the future include Segue 1 [48], Omega Centauri [49], or the newly-discovered Ursa Major III [50]. The correction happens to be particularly prominent for Sgr-like systems due to both their large spatial extension and the proximity to the Galactic center. Quantifying this correction factor for subhalos of different size and at different distances will be a topic for future analysis.

### v2

@@ 5 Gamma-ray data and model fitting
We consider a $15^{\circ}\times 15^{\circ}$ ROI centered on M54/Sgr. For our likelihood maximization, we take a 0.1 angular pixelation and use the MINUIT optimizer method within gtlike. We use an input source model that includes all sources in the 4FGL-DR3 catalog [12] out to a square of $20^{\circ}\times 20^{\circ}$. Including sources beyond the ROI ensures that sources on the edge of the ROI are properly modeled. For the interstellar emission model we use the recommended gll_iem_v07.fits, and for the isotropic emission we use iso_P8R3_SOURCE_V3_V1.txt. To report the significance of our sources, we report the TS value obtained upon running gta.sed, which takes the source following the procedure outlined above and then performs a bin-by-bin likelihood fit assuming a power law spectralization per bin with an index set to $\Gamma=-2$. Finally, for all TS values reported we assume a DM annihilation spectrum, as defined by DMFitFunction provided in FermiPy assuming annihilation via the $b\bar{b}$ channel. For this spectrum, we fit the $\mathcal{J}$-factor to the total integrated $\mathcal{J}$-factor for a given source map, fix the mass to the best-fitting mass found in ref. [13] of 29.6 GeV, and allow $\langle\sigma v\rangle$ to float.
@@
The flux of annihilating DM from the MW halo is larger than that of Sgr for s-wave, and dominates the Sgr source for p and d-waves. We obtain $\mathcal{J}_{\text{Sgr}}/\mathcal{J}_{\text{MW}}\sim 0.1-1$ for s-wave, $\mathcal{J}_{\text{Sgr}}/\mathcal{J}_{\text{MW}}\sim 0.003-0.03$ for p-wave, and $\mathcal{J}_{\text{Sgr}}/\mathcal{J}_{\text{MW}}\sim 0.001-0.005$ for d-wave, where $\mathcal{J}_{\text{MW}}$ denotes the MW $\mathcal{J}$-factor integrated over the area of the target. For the Sommerfeld case, the Sgr source flux is always larger than the MW halo emission, with $\mathcal{J}_{\text{Sgr}}/\mathcal{J}_{\text{MW}}\sim 2-7$. When fitting both the Sgr source and the MW diffuse over the entire ROI, we find an overall poor fit quality, because we are modeling the diffuse emission over such a large region of sky. This likely means that an appropriate handling of the MW DM emission requires unique templates for the diffuse MW emission arising from cosmic ray interactions with gas and the interstellar radiation field, which is beyond the scope of this work. Therefore, for the analysis in this section we model only the emission associated with the Sgr source. In this way, we can compare qualities of the source independent of the MW diffuse emission, such as the effect of including the unbound DM particles in our templates.
@@
To model M54/Sgr in gamma-rays based on our simulations results, using the results from figure 5, we create a flux map from our pixel-by-pixel calculation of the $\mathcal{J}$-factor. Since the simulated Sgr source is not smooth, we smooth the counts maps obtained from the simulation with a Gaussian kernel of size 0.2, ensuring that the total $\mathcal{J}$-factor for each map is preserved. We find that this smoothing step is necessary to obtain a good model of the source, which is generally best fit by a point source model, or a smooth model with a small, $\lesssim$ 1 degree extension [13]. Since Sommerfeld sources are the most promising candidates for detection, we focus on these models for the remainder of this section. For this portion of the analysis, we focus on the results of the Sgr analogue with the smallest angular size, Sgr 25 (angular size $\sim 1.4^{\circ}$), and compute TS values for the Sommerfeld model. Choosing the Sgr analogue with the smallest angular size ensures the best fit to the source given its small extension, as attempting to fit to the other larger analogues produces poor overall fit qualities within the ROI. For the Sommerfeld model of Sgr 25, we find a TS $\sim 7$, which corresponds to only $\sim 2\sigma$ preference for the extended source model. This includes both bound and unbound DM particles, as described above. For a map which includes only the bound DM particles to the Sgr analogue, we again find a TS of $\sim 7$.
@@ 6 Discussion and conclusions
Using analogues of the Sgr dSph identified in the Auriga simulations, we have computed $\mathcal{J}$-factors for general models in which the annihilation cross section is velocity-dependent. We provide predictions for both the DM component that is associated with the Sgr dSph, and with the MW foreground component along the line-of-sight to the dSph. For the Sommerfeld model, we find that emission from the central source is generally larger than that from the foreground MW component in the central region ($\lesssim 0.5^{\circ}$) and therefore would most likely be identifiable as a point source in the data. We further find that the $\mathcal{J}$-factors that we compute from the simulations are in agreement with those obtained from equilibrium models presented in ref. [2], implying that the latter are accurate for the central core region of Sgr.
@@
For p and d-wave models, for the first time we establish the existence of an enhancement factor due to unbound DM particles from the MW that are spatially overlapping with Sgr. These unbound particles have a relative velocity distribution characteristic of the smooth DM component of the MW at the radius of the Sgr analogue. The enhancement factor results from annihilation of these unbound DM particles with themselves and with DM particles bound to the Sgr analogues. For p-wave models, we find that this enhancement varies from $\sim 2-30$, depending on the exact Sgr analogue used. For d-wave models, we find a larger enhancement varying between $\sim 20-4000$. This enhancement is different than the more traditional “boost” factor that has been discussed often in the literature, which results from DM substructure [3, 47, 48]. On the other hand, for Sommerfeld models, the enhancement is minimal since the unbound DM particles modify the relative velocity distribution at high speeds, which is suppressed due to the $1/v$ dependence of the Sommerfeld model. For the s-wave models, the enhancement is also minimal, due to the low density of the unbound DM particles compared to the bound DM particles.
@@
Due to the angular extension of our models of $\sim 1-3$ degrees, we find that our Sgr templates generally do not provide as good a fit to the gamma-ray data as a simpler point source model. For example, examining in detail the Sommerfeld model for the case of our Sgr analogue with the smallest angular extent (Sgr 25, with an angular extent of 1.4), on Fermi-LAT data, we find a TS $\sim 7$ for the template including emission from bound and unbound particles. The TS remains the same when restricting to only the particles identified as bound to the Sgr analogue. We note that this TS value is less significant than the TS values for the point-like scenario for the M54/Sgr source as reported in refs. [12] and [13]. More detailed modeling of combined emission from the MW halo and Sgr analogue DM annihilation would require specialized detailed diffuse model templates for the region, which is beyond the scope of this work.
An effect not considered in this work is the presence of the Large Magellanic Cloud (LMC), which could have a substantial impact on the disruption of MW dwarf galaxies [34]. It is in general difficult to find exact LMC analogues in cosmological simulations. In particular, in the Auriga simulations, LMC analogues are identified only at higher redshifts and within the last 8 Gyrs, but not at the present day [49]. However, since in this work we match the properties of the Sgr analogue to its observed properties, we do not expect the existence of an LMC analogue in the simulations to significantly impact our results.
@@
We finally note that it is likely that the enhancement factor that we derive is not restricted to just Sgr-like systems. In fact, it should exist for all subhalos that are within close enough proximity to the Galactic center where there is a significant DM density that is spatially coincident with the subhalo or for subhalos that are in an overdense region in the Galaxy. Furthermore, a large difference between the average relative velocity of the DM particles bound to the subhalo and to the smooth MW halo in that region also contributes to this enhancement factor. Particularly interesting systems to investigate in the future include Segue 1 [50], Omega Centauri [51], or the newly-discovered Ursa Major III [52]. The enhancement happens to be particularly prominent for Sgr-like systems due to both their large spatial extension and the proximity to the Galactic center. Quantifying this enhancement factor for subhalos of different size and at different distances will be a topic for future analysis.